OpenStudy (anonymous):

Solve the equation. Round the answer to the nearest tenth. 2.8x = 5.1 A. 1.7 B. 1.8 C. 1.9 D. 18.2

OpenStudy (anonymous):

so you have 2.8 * x = 5.1

OpenStudy (anonymous):

how would you get rid of 2.8 to have only x on one side ?/ I.e x= ??

OpenStudy (anonymous):

what are you talking about i am confused and not that smart

OpenStudy (anonymous):

2.8x = 5.1 means that if you multiply 2.8 by an unknown number "x" it equals to 5.1

OpenStudy (anonymous):

if you have Left Hand Side (LHS) = Righ Hand Side (RHS) anything you do to one side you also do to the other.

OpenStudy (anonymous):

oo ok i get it now

OpenStudy (anonymous):

did you get an answer?

OpenStudy (anonymous):

yes i got 1.9

OpenStudy (anonymous):

i have a few more plz help me more

OpenStudy (anonymous):

Solve the equation. Round the answer to the nearest tenth. 1.26x + 2.16 = 7.71 A. 4.3 B. 9.6 C. 4.5 D. 4.4

OpenStudy (anonymous):

i got a is that right

OpenStudy (anonymous):

try recalculating

OpenStudy (anonymous):

what's yor first step

OpenStudy (anonymous):

i did 1.26 x 4.3 + 2.16 and i got 7.578 and rounded to the nearest 10th

OpenStudy (anonymous):

is that right

OpenStudy (anonymous):

plz help

OpenStudy (anonymous):

Help in what ?:)

OpenStudy (anonymous):

Solve the equation. Round the answer to the nearest tenth. 1.26x + 2.16 = 7.71 A. 4.3 B. 9.6 C. 4.5 D. 4.4

OpenStudy (anonymous):

Change the place of 2.16 to have numbers together ! Can you do this ?:)

OpenStudy (anonymous):

@86camaro

OpenStudy (anonymous):

i dont understand

OpenStudy (anonymous):

In all of the equations if we want to get answer should have numbers together and Unknowns together ! So ... Now :  Got it ?:)